# ScarFinder: a detector of optimal scar trajectories in quantum many-body dynamics

### v1

@@ A.2 Type-II scar tower
This term, however, still contains certain structure. To get a fully thermalizing perturbation, we use the projective embedding method [63] on 3-site cluster:
@@
where $\hat{P}^{[3]}$ is the local projector.
@@ Appendix C Matrix-product-state scar tower for spin-1 XY model
Figure 9: (a)-(b) Convergence behavior of ScarFinder initialized from random states with energy correction (at target energy $E_{\text{target}}=0$), with $\Delta t=0.5$. (a) For the Hamiltonian with $\hat{V}=\hat{V}_{2}$, with the perturbation in Eq. (8), among 200 independent trials, only 2 successfully converge. (b) For the modified Hamiltonian with $\hat{V}=\hat{V}_{2}^{\prime}$, with the perturbation in Eq. (15), 18 out of 20 trials converge successfully. (c)-(d) Bipartite entanglement entropy of eigenstates for the spin-1 XY Hamiltonian [Eq. (7)] with (c) $\hat{V}=\hat{V}_{2}$ and (d) $\hat{V}=\hat{V}_{2}^{\prime}$. The results are obtained via exact diagonalization on a finite system of size $L=10$, combining the momentum sectors $k=0$ and $k=\pi$.
@@
As discussed in Appendix A, besides $\hat{V}_{2}$, we can construct a more generic perturbation $\hat{V}_{2}^{\prime}$ [Eq. (26)]. Fig. 9(a)-(b) compare the convergence behaviour of ScarFinder with these two types of perturbations. Fig. 9(c)-(d) contrast the entanglement entropy spectra for these two perturbations, $\hat{V}_{2}$ and $\hat{V}_{2}^{\prime}$. Compared to panel (c), panel (d) exhibits a clearer separation between low-entanglement scar states and highly entangled thermal states. This sharper distinction suggests that the modified perturbation $\hat{V}_{2}^{\prime}$ creates a more favorable spectral structure and improves the performance of the ScarFinder algorithm.
Appendix D Mixed field Ising model
@@
Figure 10: Eigenstate overlaps for optimized initial states obtained via ScarFinder at various target energy densities $\epsilon$ in the mixed-field Ising model, Eq. (30). Each initial state is optimized within the iMPS manifold of bond dimension $\chi=10$. The optimal states are selected from an ensemble of 100 random initializations, with a time step $\Delta t=0.5$ over 1000 iterations. The optimal state at each energy density is chosen as the one exhibiting minimal entanglement entropy at $t=4.0$. The overlap data are obtained via exact diagonalization of the Hamiltonian Eq. 30 (with $(J=1.0,h=0.5,g=1.05$) on a system of size $L=16$, restricted to the momentum sector $k=0$, which corresponds to two-site translational symmetry. This symmetry reduction arises from the two-site unit cell used in the iTEBD simulation, which explicitly breaks the original one-site translation symmetry down to two-site periodicity.
@@
The transverse-field Ising model with an additional longitudinal field provides a canonical example of a nonintegrable quantum many-body system:
@@
For such nonintegrable quantum Hamiltonians, a generic initial state is expected to rapidly thermalize: local observables should swiftly equilibrate to their thermal expectation values, and the bipartite entanglement entropy should quickly saturate to a volume-law scaling. However, a previous study [74] has revealed exceptions to this expectation, where in the chaotic regime (parameters set as $J=1.0$, $h=0.5$, $g=1.05$), certain initial product states, such as the fully polarized state $|Z+\rangle\equiv\left|\uparrow\uparrow\cdots\uparrow\right\rangle$, exhibit anomalously slow thermalization dynamics. This phenomenon, referred to as “weak thermalization”, was subsequently attributed to high overlaps of the initial state with low-entanglement eigenstates in the vicinity of the system’s ground state [75].
Motivated by these observations, we apply the ScarFinder algorithm to systematically search for optimized scar-like initial states in this nonintegrable model. In contrast to our previous investigations of the PXP model in the main text, where it was sufficient to set the targer energy to zero and study the middle of the spectrum, for the mixed-field Ising model in Eq. (30) we expect the results to be more sensitive to the energy density. Hence, we vary the target energy density: $\epsilon=(E-E_{0})/(E_{\text{max}}-E_{0})$, with $E_{0}$ and $E_{\text{max}}$ denote the lowest and highest eigenenergies, respectively. Figure 10 shows the eigenstate overlaps of optimized initial states identified via ScarFinder at different energy densities, demonstrating distinctive spectral signatures, especially at higher energy densities.
Figure 11: Entanglement entropy growth comparison between the optimized ScarFinder initial state and the product state in Eq. (31) with $|\theta\approx 0.1649\pi\rangle$, both at energy density $\epsilon=0.1$. The optimized initial state demonstrates significantly suppressed entanglement growth compared to the polarized state.
To illustrate the improvement provided by ScarFinder, we compare the dynamics of optimized states against simple product states from Ref. [74], when the two are matched to have the same energy density. Specifically, we choose the product state:
@@
with $\theta\approx 0.1649\pi$, yielding an energy density $\epsilon=0.1$. Figure 11 compares the entanglement entropy dynamics from this product state against the ScarFinder’s optimized initial state. The optimized state exhibits considerably slower entanglement growth, demonstrating the algorithm’s efficacy in isolating dynamically atypical initial conditions with suppressed thermalization.
@@
While this demonstrates that the ScarFinder algorithm can successfully identify nonergodic dynamics in generic nonintegrable models, further work is needed to fully understand the underlying mechanisms of these emergent trajectories. In particular, in the intermediate energy density range, e.g., $\epsilon=0.4$ or $\epsilon=0.8$ in Fig. 10, one notices intriguing tower structures that are somewhat reminiscent of the energy-dependent QMBSs in the PXP model in the presence of a chemical potential [23, 72]. It would be interesting to understand if these are indeed related to QMBSs or if the quasiparticle picture from Ref. [75] can be generalized to explain these atypical spectral features.

### v3

@@ A.2 Type-II scar tower
Nevertheless, this term still contains some structure. To get a fully thermalizing perturbation, we use the projective embedding method [ShiraishiMori] on a 3-site cluster:
@@
where $\hat{P}^{[3]}$ is the local projector such that $(1-\hat{P}^{[3]})$ locally annihilates all scar states.
@@ Appendix C Matrix-product-state scar tower for spin-1 XY model
Figure 9: (a)-(b) Convergence behavior of ScarFinder for the spin-1 XY model initialized from random states with energy correction (at target energy $E_{\text{target}}=0$), with $\Delta t=0.5$. (a) For the Hamiltonian with $\hat{V}=\hat{V}_{2}$ in Eq. (25), among 200 independent trials, only 2 successfully converge. (b) For the Hamiltonian with $\hat{V}=\hat{V}_{2}^{\prime}$ in Eq. (26), 18 out of 20 trials converge successfully. (c)-(d) Bipartite entanglement entropy of eigenstates for the spin-1 XY Hamiltonian [Eq. (7)] with (c) $\hat{V}=\hat{V}_{2}$ and (d) $\hat{V}=\hat{V}_{2}^{\prime}$. The results are obtained via exact diagonalization on a finite system of size $L=10$, combining the momentum sectors $k=0$ and $k=\pi$.
@@
As discussed in Appendix A, besides $\hat{V}_{2}$, we can construct a more generic perturbation $\hat{V}_{2}^{\prime}$ [Eq. (26)]. Fig. 9(a)-(b) compare the convergence behaviour of ScarFinder with these two types of perturbations. Fig. 9(c)-(d) contrast the entanglement entropy spectra for these two perturbations, $\hat{V}_{2}$ and $\hat{V}_{2}^{\prime}$. Compared to panel (c), panel (d) exhibits a clearer separation between low-entanglement scar states and highly-entangled thermal states. This sharper distinction suggests that the modified perturbation $\hat{V}_{2}^{\prime}$ creates a more favorable spectral structure that improves the performance of the ScarFinder algorithm.
Appendix D PXP-type models on quasi-1D lattices
While the main text focuses on 1D models, the core ideas behind ScarFinder are not restricted to that setting. In principle, the algorithm can be extended to higher-dimensional systems, provided one can define a suitable low-entanglement variational manifold. However, such generalizations face significant technical challenges, as entanglement structures are more complex in higher dimensions and tensor network methods are much more difficult to implement [CiracRMP]. As a practical intermediate step, we explore quasi-1D lattice geometries which go beyond simple chains in structure, but retain limited width, allowing for efficient simulation using MPS-based methods. In the following, we apply ScarFinder to the PXP model defined on square and triangular lattice tubes, illustrated in Fig. 10. We demonstrate that the algorithm remains effective in uncovering scarred dynamics, while it also allows to find exact eigenstates even in these extended quasi-1D settings.
Figure 10: Quasi-1D versions of the PXP model defined on (a) square lattice and (b) triangular lattice. Both systems form cylindrical tubes with periodic boundary conditions along the vertical direction and infinite length horizontally. The Rydberg blockade constraint is imposed along the nearest-neighbor links. Each block, denoted by dashed line, outlines a unit cell that is repeated along the horizontal axis. The unit cell is composed of four sublattice sites labeled $\mathrm{A},\mathrm{B},\mathrm{C},\mathrm{D}$, which are used to define the MPS ansatz for ScarFinder.
Figure 11: Dynamics of the PXP model on a quasi-1D lattice. (a) Entanglement entropy evolution starting from the charge density wave (CDW) state and from optimized initial states obtained by ScarFinder with various bond dimensions $\chi=2,4,8$. The optimized states exhibit suppressed entanglement growth compared to the CDW state. (b) Local spin expectation values $\langle\hat{Z}\rangle$ on two sublattices show nearly perfect revival dynamics from the optimized state with $\chi=8$. We ran ScarFinder with projection time step $\Delta t=1.0$ for 200 iterations and the underlying dynamics were simulated using the iTEBD algorithm with a maximum bond dimension of $\chi_{\text{max}}=24$.
D.1 Square lattice on a thin cylinder
Previously, for the PXP model on a 2D square lattice, Refs. [lin2020quantum, PhysRevResearch.2.022065] studied the dynamics from a charge-demsity-wave state, $|\text{CDW}\rangle$, with excitations on every site of one of the two sublattices of this bipartite lattice. Qualitatively, the scar dynamics were found to behave similarly as in the 1D case. As a first extension, we apply ScarFinder to the PXP model defined on a square lattice, but wrapped onto a finite cylinder shown in Fig. 10(a). We consider a cylinder defined by periodic boundary conditions along one spatial dimension, while leaving the other dimension infinite. In this quasi-1D setting, the lattice can be described using MPS methods by grouping sites around the circumference into composite sites, as depicted by dashed lines in Fig. 10(a).
Specifically, we study the square lattice with circumference $L_{y}=4$ and infinite length along the $x$-direction. We apply ScarFinder by employing translationally-invariant MPS representation adapted to this geometry, where each physical site represents one circumference unit cell of the cylinder. Note that, because of the constraint, the local Hilbert space dimension for each blocked site is $d=7$. For notational simplicity, we define the following 4-site configurations of the block:
[…]
The CDW state can be naturally expressed as a product of alternating $|\mathbb{Z}_{2}\rangle$ and $|\mathbb{Z}_{2}^{\prime}\rangle$ configurations: $|\text{CDW}\rangle=|\mathbb{Z}_{2},\mathbb{Z}_{2}^{\prime},\dots\rangle$.
Now we can apply ScarFinder starting from the $|\text{CDW}\rangle$ state and choosing bond dimensions $\chi=2,4,8$ to find an optimized scar initial state. The resulting entanglement dynamics are shown in Fig. 11(a). Compared to the CDW state, which exhibits steadily increasing entanglement the optimized states found by ScarFinder display slower entanglement growth with clearly visible revivals, though sharing the same revival frequency as the $|\text{CDW}\rangle$ state. As the bond dimension increases from $\chi=2$ to $\chi=8$, these revivals become more pronounced, and the entanglement entropies become substantially lower than that of the $|\text{CDW}\rangle$ state at late times.
To further illustrate the coherent dynamics, Fig. 11(b) shows the time evolution of local observables $\langle\hat{Z}_{1}\rangle$ and $\langle\hat{Z}_{2}\rangle$, representing magnetization on two underlying sublattices. The data correspond to dynamics initialized from the optimized state with $\chi=8$ and reveal stable periodic oscillations. These oscillations qualitatively resemble those of the $|\text{CDW}\rangle$ state, but the optimization provided by ScarFinder yields notably improved revival stability and reduced entanglement growth. These results confirm that ScarFinder effectively enhances the scar dynamics of the quasi-1D PXP model defined on a finite cylinder. While numerical limitations currently restrict our analysis to such quasi-1D settings, it would be interesting to verify if a similar optimization procedure could also yield improved scar initial states in the full 2D PXP model.
D.2 Triangular lattice on a thin cylinder
As a further application, we apply ScarFinder to the PXP model on a triangular lattice wrapped around a cylinder, with the vertical circumference of four lattice sites and infinite length in the horizontal direction [see Fig. 10(b)]. QMBSs on non-bipartite lattices, such as the triangular lattice, have remained largely unexplored and no scar states have previously been reported in such settings, including the experimental study in Ref. [bluvstein2021controlling]. Below we demonstrate that the ScarFinder successfully identifies an exact scar eigenstate for the geometry in Fig. 10(b).
Specifically, we perform random sampling over initial states and apply ScarFinder using a projection time step of $\Delta t=1.0$ for 100 iterations. For bond dimensions $\chi>2$, the algorithm consistently converges to a unique low-entanglement initial state. Upon inspection, we find this state can be exactly represented with $\chi=2$. Numerical checks confirm that this identified state remains invariant under the PXP Hamiltonian dynamics, revealing itself as an exact eigenstate.
To analyze the structure of the found MPS state, we first note that MPS tensors have nonzero entries only when the physical indices correspond to the three states in Eq. (30). Thus, the local Hilbert space is effectively reduced to a spin-1 degree of freedom per site. To further characterize the state, we restrict the system to a finite periodic chain of length $L=4$, convert the MPS into a state vector, and inspect the nonzero amplitudes. We find that these nonzero coefficients all have magnitude $1/\sqrt{17}$ and differ only by sign. The nonzero amplitudes correspond exactly to product-state configurations that avoid placing two adjacent $|\mathbb{Z}_{2}\rangle$ or $|\mathbb{Z}_{2}$$\rangle$ blocks. For a length-four periodic system, precisely 17 such allowed configurations exist. The sign of each amplitude depends on the parity (even or odd positions) of the $|\mathbb{Z}_{2}\rangle$ and $|\mathbb{Z}_{2}$$\rangle$ states. Based on these insights, we determine the exact wave function
$|S\rangle=\mathcal{P}\bigotimes_{j}\left(|0\rangle_{j}-|\mathbb{Z}_{2}\rangle_{j}-|\mathbb{Z}_{2}’\rangle_{j}\right),$ (31)
where the projector $\mathcal{P}$ eliminates any configurations containing adjacent $|\mathbb{Z}_{2}\rangle$ or $|\mathbb{Z}_{2}$$\rangle$ blocks.
Given this form, one can straightforwardly verify that the state is annihilated by the PXP Hamiltonian. In particular, each local term $\hat{h}_{j}=\mathcal{P}\hat{X}_{j}\mathcal{P}$, where $\hat{X}_{j}=\hat{\sigma}_{j,A}^{x}+\hat{\sigma}_{j,B}^{x}+\hat{\sigma}_{j,C}^{x}+\hat{\sigma}_{j,D}^{x}$ is the sum of Pauli-$X$ operators acting on all sublattice spins at site $j$, individually annihilates the state $|S\rangle$, as shown below:
$\hat{h}_{j}|S\rangle=$ $\ \mathcal{P}\!\left[\cdots\otimes\hat{h}_{j}|0\rangle_{j-1}\otimes(|0\rangle_{j}-|\mathbb{Z}_{j}\rangle\right$. (32)
$\ \left.-|\mathbb{Z}^{\prime}_{j}\rangle)\otimes|0\rangle_{j+1}\otimes\cdots\right]$
$=$ $\ \mathcal{P}\!\left[\cdots\otimes|0\rangle_{j-1}\otimes\hat{X}_{j}(|0\rangle_{j}-|\mathbb{Z}_{j}\rangle\right$.
$\ \left.-|\mathbb{Z}^{\prime}_{j}\rangle)\otimes|0\rangle_{j+1}\otimes\cdots\right].$
The action of $\hat{X}$ on site $j$ gives
$\hat{X}(|0\rangle-|\mathbb{Z}\rangle-|\mathbb{Z}^{\prime}\rangle)$ (33)
[…]
Each of these configurations contains adjacent excitations and is therefore projected out by $\mathcal{P}$, which enforces the Rydberg blockade constraint. Consequently, $\hat{h}_{j}|S\rangle=0$ for every $j$, implying that the full PXP Hamiltonian annihilates the state:
$\hat{H}_{\mathrm{PXP}}|S\rangle=0.$ (34)
Thus, $|S\rangle$ is an exact eigenstate with eigenvalue $E=0$.
We emphasize that the exact cancellations discussed above occur as a special case arising from the $4\times\infty$ cylindrical geometry. It would be interesting to understand if some generalization of this construction may apply to quasi-1D geometries with a larger number of horizontal layers or even in the isotropic 2D limit. Furthermore, we note that our search on a $4\times\infty$ cylinder did not reveal any long-lived dynamical scar states (beyond the exact eigenstate mentioned above), suggesting that robust scar dynamics may not be present in more general non-bipartite systems such as the 2D triangular lattice.
Appendix E Mixed field Ising model
@@
Figure 12: Eigenstate overlaps for optimized initial states obtained via ScarFinder at various target energy densities $\epsilon$ in the mixed-field Ising model, Eq. (35). Each initial state is optimized within the iMPS manifold of bond dimension $\chi=10$. The optimal states are selected from an ensemble of 100 random initializations, with a time step $\Delta t=0.5$ over 1000 iterations. The optimal state at each energy density is chosen as the one exhibiting minimal entanglement entropy at $t=4.0$. The overlap data are obtained via exact diagonalization of the Hamiltonian Eq. 35 (with $(J=1.0,h=0.5,g=1.05$) on a system of size $L=16$, restricted to the momentum sector $k=0$, under two-site translational symmetry. This symmetry reduction arises from the two-site unit cell used in the iTEBD simulation, which explicitly breaks the original one-site translation symmetry down to two-site periodicity.
@@
The transverse-field Ising model with an additional longitudinal field is a canonical example of a nonintegrable quantum many-body system:
@@
For such Hamiltonians, a generic initial state is expected to rapidly thermalize: local observables should swiftly equilibrate to their thermal expectation values, and the bipartite entanglement entropy should quickly saturate to a volume-law scaling. However, previous studies [Banuls2011, Alishahiha2025] revealed exceptions to this expectation, where in the chaotic regime (parameters set as $J=1.0$, $h=0.5$, $g=1.05$), certain initial product states, such as the fully polarized state $|Z+\rangle\equiv\left|\uparrow\uparrow\cdots\uparrow\right\rangle$, exhibit anomalously slow thermalization dynamics. This phenomenon, referred to as “weak thermalization”, was subsequently attributed to high overlaps of the initial state with low-entanglement eigenstates in the vicinity of the system’s ground state [LinMotrunich2017]. Recent works, however, have raised questions about whether the model thermalizes in the thermodynamic limit [PhysRevB.106.214311], or whether thermal behavior only emerges at much larger system sizes than previously explored [arxiv:2409.18863].
Motivated by these observations, we apply the ScarFinder algorithm to systematically search for optimized scar-like initial states in this nonintegrable model. In contrast to our previous investigations of the PXP model in the main text, where it was sufficient to set the targer energy to zero and study the middle of the spectrum, for the mixed-field Ising model in Eq. (35) we expect the results to be more sensitive to the energy density. Hence, we vary the target energy density $\epsilon=(E-E_{0})/(E_{\text{max}}-E_{0})$, with $E_{0}$ and $E_{\text{max}}$ denoting the lowest and highest eigenenergies, respectively. Fig. 12 shows the eigenstate overlaps of optimized initial states identified via ScarFinder at different energy densities, demonstrating distinctive spectral signatures, especially at higher energy densities.
Figure 13: Comparison of the optimized ScarFinder initial state and the polarized product state $|\theta\approx 0.1649\pi\rangle$ from Eq. (36), both at energy density $\epsilon=0.1$. (a) Entanglement entropy growth comparison between the optimized ScarFinder initial state and the product state in Eq. (36). The optimized initial state demonstrates significantly suppressed entanglement growth compared to the polarized state. (b) Logarithmic fidelity $\mathcal{N}[\psi_{t},\psi_{0}]$ as a function of time. The optimized state shows larger and more regular revivals, further confirming its enhanced coherence.
To illustrate the improvement provided by ScarFinder, we compare the dynamics of optimized states against simple product states from Ref. [Banuls2011], when the two have the same energy density. Specifically, we choose the product state:
@@
with $\theta\approx 0.1649\pi$, yielding an energy density $\epsilon=0.1$. Fig. 13(a) compares the entanglement entropy dynamics from this product state against ScarFinder’s optimized initial state. The optimized state exhibits considerably slower entanglement growth, demonstrating the algorithm’s efficacy in isolating dynamically atypical initial conditions with suppressed thermalization. In addition, Fig. 13(b) shows the logarithmic fidelity dynamics, where the optimized state maintains larger and more regular revivals compared to the polarized state, further supporting its nonthermal character.
@@
While this demonstrates that the ScarFinder algorithm can successfully identify nonergodic dynamics in generic nonintegrable models, further work is needed to fully understand the underlying mechanisms of these emergent trajectories. In particular, in the intermediate energy density range, e.g., $\epsilon=0.4$ or $\epsilon=0.8$ in Fig. 12, one notices intriguing tower structures that are somewhat reminiscent of the energy-dependent QMBSs in the PXP model in the presence of a chemical potential [Su2023, Daniel2023]. It would be interesting to understand if these are indeed related to QMBSs or if the quasiparticle picture from Ref. [LinMotrunich2017] can be generalized to explain these atypical spectral features.